OSMOS Mechanism Numbering

Inside the IEB

Here is a schematic of the inside of the instrument electronics box with the MicroLYNX numbered.

Outside the IEB

Here is how the numbering corresponds to the mechanism cable positions on the IEB panel. These numbers correspond to the individual MicroLYNX. Mechanism cables can be swapped as necessary to the open position if a MicroLYNX fails. The code for that mechanism must be uploaded into the spare MicroLYNX.

Outside OSMOS

Here is how the positions on the OSMOS panel correspond to the mechanisms within the instrument. This does not change.

Original Configuration

The as-shipped configuration was as follows:

ML# Port Mechanism ML Code
18001Slit Wheelslit
28002Collimator Focuscolfoc
38003Disperser Wheeldisperser
48004Camera Focuscamfoc
58005Filter Wheel 1filter1
78007Filter Wheel 2filter2

Current Configuration

The current configuration (as of 2017 Aug 24) is:

ML# Port Mechanism ML Code Date By
18001Slit Wheelslit2015 Feb 5rwp/osu
28002Collimator Focuscolfoc2014 Feb 7rwp/osu
38003deadn/a2015 Dec 28rwp/osu
48004Camera Focuscamfoc2012 Octrwp/osu
58005Disperserdisperser2015 Oct 13rwp/osu
68006Filter Wheel 1filter12015 Feb 5rwp/osu
78007Filter Wheel 2filter22014 Feb 7rwp/osu

Trouble Record by MicroLYNX position

2012 Oct: Replaced with new MicroLYNX Oct 2012, started showing intermittent failures in May 2012, but no outright failures or the red fault LED light.

2014 Feb: replaced unit, original showing flaky behavior.

2015 Dec 28: Failed, showing red LED indicator, swapped disperser to MLC5
2012 Oct: showing flaky behavior, replaced with new unit

2012 Oct: Failed (red LED), replaced with new unit.

2015 Feb 5: Replaced failed unit

2015 Feb 5: Replaced Failed unit.

2014 Feb: Replaced after orginal failed Jan 2014
The large number of failures and flaky behavior in October 2012 followed a thunderstorm with frequent cloud-to-ground lightning near the observatory, so the probably cause was lightning surge damage.

Additional surge protection circuitry was introduced in 2015 following troubleshooting of failed/flaky units to OSU for testing.

Updated: 2017 Aug 24 [rwp/osu]